# Towards Quantifying Long-Range Interactions in Graph Machine Learning: a Large Graph Dataset and a Measurement

### v1

@@ 3 Benchmarks: From Classical GNNs to Graph Transformers
In this section, we benchmark our dataset by comparing the empirical performance of classical GNNs and Graph Transformers under different experimental settings. We emphasize at the outset that the primary objective of our experiments is to demonstrate the presence of long-range dependencies in the proposed City-Networks with associated task, rather than maximizing each model’s performance with its optimal hyperparameters. Readers are referred to Appendix B for more training details and our experiment settings.
Figure 3: Results for setting number of layers $L$ to match number of hops $H\in[2,4,8,16]$.
Experimental setups.
We consider transductive node classification and split train/validation/test sets by $10\%/10\%/80\%$ on all networks, where we fix the data split for benchmarking purposes, and adopt the average classification accuracy as our evaluation metric. For baselines, we consider both spectral and spatial GNNs: ChebNet (Defferrard et al., 2016), GCN (Kipf & Welling, 2017), and GraphSAGE (Hamilton et al., 2017); and a scalable Graph Transformer, SGFormer (Wu et al., 2023), which combines local message-passing and linear global attention. In addition, we also include an MLP without using graph information to reflect the importance of topology in our task. All the baselines are implemented with standard methods in PyTorch Geometric (PyG) (Fey & Lenssen, 2019) if available or otherwise based on their officially released codes. During training, all models are optimized for $20k$ epochs using AdamW (Loshchilov & Hutter, 2019) with a learning rate of $10^{-3}$ and a weight decay of $10^{-5}$. At the end of training, we select the epoch with the best validation performance and record its corresponding test accuracy. We then report the mean and standard deviation of the test accuracies across 5 runs.
Scalable training with $H$-hop sampling.
Since the full-batch training is infeasible on our large-scale city networks, we adopt a standard mini-batching method (Hamilton et al., 2017) implemented by PyG – NeighborLoader, which first samples $S$ seed nodes and grows an $H$-hop neighborhood for each seed node (i.e. an $H$-hop ego-network). At each iteration in the loader, these subgraphs will then be batched together and sent to the model to generate predictions on the seed nodes only. Importantly, this method implies that the model can only observe information from maximally $H$ hops, which enables us to investigate the long-range dependency by controlling $H$ and then comparing the model’s performance. Specifically, we test each baseline with different numbers of hops at $H\in[2,4,8,16]$ under two settings: (1) matching the number of layers $L$ to $H$, and (2) fixing the number of layers to $L=16$ (i.e. freeze the model size) for different $H$. For experimental purposes, we employ a fixed hidden channel size of $32$ in all baselines to maintain a comparable parameter size across models.
Results.
The results for settings (1) and (2) are shown in Figure 3 and Figure 4, respectively. When matching $L=H$ and gradually increasing the number of sampled hops $H$, we can observe a clear improvement in performance for all baselines on our four city networks, while on Cora, the baselines all suffer from over-smoothing, except for SGFormer  [^3] We use the default split on Cora for all baselines rather than the random split from the SGFormer paper., whose performance remains relatively stable. This result suggests that incorporating information from more distant hops will improve both GNNs and Graph Transformers’ performance on City-Networks, but such gain is outweighed by over-smoothing issues on Cora due to the increasing number of GNN layers. On the other hand, one could argue that the observed performance gains may be attributed to the increasing model parameters as the number of layers $L$ grows. To address this, we further investigate the scenario where $L$ is fixed at 16 (i.e., keeping the model size constant) while gradually increasing $H$. Similar to the previous results, all baselines exhibit a consistent upward trend on our proposed dataset, whereas no clear pattern emerges on Cora (all baselines are subject to over-smoothing in this case). We interpret these results that long-range dependency is the primary factor that contributes to the improvement of baselines’ performance as $H$ increases, and we proceed to introduce a quantitative measure of such dependency in the next section.
Table 2: Baseline results of the best testing accuracy on City-Networks.
Baseline Paris Shanghai Los Angeles London
MLP $23.6\pm 0.8$ $27.5\pm 0.3$ $23.1\pm 0.6$ $27.0\pm 0.3$
ChebNet $49.5\pm 0.4$ $57.9\pm 1.4$ $54.1\pm 0.2$ $49.4\pm 0.4$
GCN $47.3\pm 0.2$ $52.4\pm 0.3$ $45.9\pm 1.0$ $43.8\pm 0.3$
GraphSAGE $49.1\pm 0.6$ $60.4\pm 0.3$ $53.3\pm 0.7$ $48.2\pm 0.8$
SGFormer $45.0\pm 0.2$ $53.5\pm 0.3$ $47.2\pm 0.5$ $45.7\pm 0.3$
For reference, we also summarize the best performance of the baselines on our City-Networks in Table 2, which, as expected, all occur at $L=H=16$. We reiterate that, through the above analysis, our primary goal is to demonstrate the long-range dependencies in the data and task manifested in model performance. It is important to note that the reported results may not reflect the optimal performance of these baselines, which constitutes a limitation of this study and is further discussed in Appendix B.
4 Quantifying Long-Rangeness: A Measurement Based on Jacobians
@@
In this section, we propose to quantify the long-range dependency of a task given a trained model by seeing how the influence score $I(v,u)$ between node-pairs  (Xu et al., 2018; Gasteiger et al., 2022) varies with distance, which measures the sensitivity of a GNN layer at node $v$ to the input feature of node $u$ using the Jacobian:
$I(v,u)=\sum_{i}\sum_{j}\absolutevalue{\partialderivative{\bm{H}^{(L)}_{vi}}{\bm{X}_{uj}}},$ (2)
where $\bm{H}^{(L)}_{vi}$ is the $i^{th}$ entry of node $v$’s embedding at layer $L$, and $\bm{X}_{uj}$ is the $j^{th}$ entry of the input feature vector for node $u$. Unless otherwise specified, we assume $L$ refers to the last logit layer (pre-softmax inputs).
@@
where $T_{h}(v)=I_{\text{sum}}(v,h)$ is the total influence from the $h^{th}$-hop neighbors of node $v$, $\rho(v,u)$ is the length of the shortest path between $v$ and $u$ (note that this is equivalent to the $h$-hop shell later discussed in Section 6.3), and $N$ is the number of nodes in the network. Also, we would like to highlight that when $h=0$, $T_{0}(v)=I(v,v)$ becomes the influence of the feature at node $v$ on its output. The average total influence quantifies how much, on average, the features of nodes that are $h$ hops away affect the output at the focal node. In other words, by summing the Jacobian-based influence scores from all $h$-hop neighbors (or shells) and then averaging over all nodes, it provides an expected measure of the cumulative effect that distant nodes have on each node as the focal node.
@@
where $H$ is the maximum number of hops to be considered. Intuitively, one can understand $R$ as measuring how far away the average unit of influence is, where graphs with long-range dependencies are expected to have higher proportional influence between more distant nodes. We refer readers to Appendix C for a discussion on the computational cost of this measurement.
@@
Table 3: The average size of the influence-weighted receptive field $R$ on different datasets, where the underlying models are trained with $L=H=16$ from Section 3.
@@
Model Paris Shanghai L.A. London Cora Citeseer ogbn-arxiv
GCN 4.01 4.58 5.34 6.19 3.25 1.53 1.33
ChebNet 6.30 4.89 6.77 7.62 2.33 0.86 2.98
GraphSAGE 5.79 5.83 6.69 7.19 2.28 0.66 1.33
SGFormer 3.86 4.62 4.05 4.65 0.77 0.51 1.53
@@
Figure 5: Normalized average total influence $\bar{T}_{h}/\bar{T}_{0}$ averaged across nodes at different distances.
Empirical results.
We validate $R$ and $\bar{T}_{h}$ on our datasets using four models from Section 3, each with $16$ layers and trained using $16$-hop neighborhood sampling (i.e., $H=L=16$). To analyze long-range dependencies, we also train these models on Cora, Citeseer, and ogbn-arxiv under identical settings and compare their behavior in this setting to the patterns on City-Networks. The results in Table 3 show that $R$ is consistently higher on our City-Networks compared to the citation networks across all models, which indicates a longer range of influence. To make $\bar{T}_{h}$ more comparable across models and datasets, we normalize it to $\bar{T}_{h}/\bar{T}_{0}$ and present the results in Figure 5. We can observe that for all models, there is a clearer and more rapid decay of influence at distant hops in all citation networks compared to city networks. We note that $R$ and $\bar{T}_{h}$ are model-dependent; however, results on these measures provide a consistent ordering between datasets: our city networks and their associated tasks require trained models to consider features from further hops compared to citation networks when generating predictions.
For the rest of this paper, we will first provide a theoretical justification of the proposed dataset by linking over-smoothing to algebraic connectivity and network diameter in Section 5, and then present the intuition behind the proposed long-range measurement in Section 6.
5 Theoretical Justification of Graph Topologies in City-Networks
@@
In this section, we provide a spectral understanding of over-smoothing in GNNs. We emphasize that, rather than offering an in-depth theoretical analysis, our aim is to justify our dataset design choices by grounding them in theoretical foundations. The section is organized as follows. First, we review the concept of information loss in the limit of infinite GNN layers due to over-smoothing. We show that over-smoothing effectively vanishes the original node features and makes representations collapse to a value that is only dependent on the graph topology. Next, we prove that the rate at which over-smoothing happens is controlled by the eigenvalues of the normalized adjacency operator in linear GCNs. Given that eigenvalues with larger absolute values slow down over-smoothing, we analyze how the magnitude of the eigenvalues relates to the graph topology of our datasets. In particular, we show that sparse graphs with large diameters promote bigger positive eigenvalues for the normalized adjacency operator and hence are less vulnerable to over-smoothing. Finally, we argue that less over-smoothing makes it possible for GNNs to build long-range representations in the case of our proposed dataset. All proofs are presented in Appendix D.1.
@@ 5.1 Preliminaries: A Spectral Perspective on Over-smoothing for Linear GCNs
We begin by analyzing over-smoothing through the lens of linearization, following the framework established in Wu et al. (2019) to build Simple Graph Convolutional Networks (SGCs). We select this model for our analysis because it is a linearized and more interpretable version of the widely adopted GCN model; in our empirical study in Section 3, we use the widely adopted classical GCN. Linear convolutions have also been used for theoretical analysis in other previous works such as Giovanni et al. (2022). Although GCNs are often regarded in recent literature as local neighborhood aggregators, they were originally derived from a spectral perspective. Thus, our analysis revisits the original conceptual foundation that motivated the model.
@@
Let $\mathcal{G}=(V,E)$ be an undirected graph with adjacency matrix $\bm{A}\in\mathbb{R}^{N\times N}$ and degree matrix $\bm{D}\in\mathbb{R}^{N\times N}$ ($\bm{D}_{ii}=\sum_{j}\bm{A}_{ij}$), where $N=|V|$ is the total number of nodes in the graph. The normalized (augmented) adjacency operator $\bm{\tilde{S}}_{adj}\in\mathbb{R}^{N\times N}$ is defined as:
$\bm{\tilde{S}}_{adj}=(\bm{I}+\bm{D})^{-\frac{1}{2}}(\bm{I}+\bm{A})(\bm{I}+\bm{D})^{-\frac{1}{2}},$ (5)
where the original adjacency matrix and degree matrix have been augmented with self-loops leveraging the identity matrix $\bm{I}$. A Graph Convolutional Network (GCN) layer with input features $\bm{X}^{(l)}$, learnable weights $\bm{W}^{(l)}$ for layer $l$, and non-linear pointwise activation function $\sigma$ is defined as:
$\bm{\tilde{S}}_{adj}=\bm{\tilde{U}}\bm{\Lambda}\bm{\tilde{U}}^{T},$ (8)
$(\bm{\tilde{S}}_{adj})^{k}=\bm{\tilde{U}}\bm{\Lambda}^{k}\bm{\tilde{U}}^{T},$ (9)
5.2 Over-smoothing Rate as a Function of Algebraic Connectivity, Diameter, and Sparsity
@@
Next, we provide intuition about the types of graph topologies that would mitigate the over-smoothing problem. In particular, sparse graphs—those lacking the small-world effect commonly found in citation networks—and graphs with large diameters tend to experience over-smoothing at a slower rate. We argue that this slower rate of over-smoothing implies a higher likelihood that GNNs can learn long-range dependencies during optimization. We show this intuition by bounding the eigenvalues of the normalized adjacency matrix using graph properties, then showing how these eigenvalues influence over-smoothing, and thus how controlling these graph properties can mitigate over-smoothing.
Recall that the normalized Laplacian matrix $\bm{L}_{sym}$ of a graph is defined as:
$\bm{L}_{sym}=\bm{I}-\bm{D}^{-\frac{1}{2}}\bm{A}\bm{D}^{-\frac{1}{2}}=\bm{I}-\bm{{S}}_{adj},$ (12)
which is intimately related to the normalized adjacency operator, $\bm{\tilde{S}}_{adj}$. Also, recall that the normalized algebraic connectivity $\lambda_{2}(\bm{L}_{sym})$ is defined as the second smallest eigenvalue of $\bm{L}_{sym}$. For connected graphs like ours, $\lambda_{2}(\bm{L}_{sym})>0$. We first present the following lemma.
Lemma 5.1 (Eigenvalue complementarity of normalized operators).
$\lambda_{N+1-i}(\bm{{S}}_{adj})=1-\lambda_{i}(\bm{L}_{sym})$ (13)
This complementarity implies that when the normalized algebraic connectivity $\lambda_{2}(\bm{L}_{sym})$ is small, the second largest positive eigenvalue of $\bm{{S}}_{adj}$ must be close to 1. However, in our context, we are interested in the algebraic connectivity of the normalized graph Laplacian that would correspond to the normalized adjacency operator, $\bm{\tilde{S}}_{adj}$, introduced in Equation (5), instead of that of $\bm{{S}}_{adj}$. Thus, we extend (Wu et al., 2019, Theorem 1) to cover the normalized algebraic connectivity:
@@
Assume a connected graph $\mathcal{G}$. For all $\gamma>0$,
@@
That is, the second largest positive eigenvalue of the normalized adjacency operator of a graph with self-loops is larger than that of the same graph without self-loops, since adding self-loops “sparsifies” the original graph and reduces its algebraic connectivity. This allows us, via Lemma 13, to relate the second largest positive eigenvalue of $\bm{\tilde{S}}_{adj}$ to the topology of graphs in the proposed dataset, using bounds similar to those presented in (Chung, 1997, Lemma 1.14).
@@
From Table 4, we see that the lower bound in Equation (15) is generally higher for our datasets; this is because the bound is decreasing in $d_{max}$ and increasing in $diam(\mathcal{G})$, showing graphs with large diameter and low maximum degree will be more resilient to over-smoothing. We provide the rationale as follows.
@@
Dataset City-Network (ours) LRGB Planetoid OGB
Paris Shanghai L.A. London PascalVOC COCO Cora CiteSeer ogbn-arxiv
Bound 0.4741 0.6344 0.6095 0.5921 0.4857 0.4815 0.0324 0.1143 $-0.0640$
The second largest eigenvalue of $\bm{\tilde{S}}_{adj}$ in terms of magnitude is either $\lambda_{N-1}$ or $\lambda_{1}$ We now explicitly consider the case when it is $\lambda_{N-1}$. As $\lambda_{N-1}$ approaches 1, the rate of convergence to the limiting state $\bm{\tilde{u}}_{N}\bm{\tilde{u}}_{N}^{T}$ decreases exponentially with the number of layers. For any layer $l$, the difference from the limiting state can be expressed as:
$||(\bm{I}-\bm{\tilde{u}}_{N}\bm{\tilde{u}}_{N}^{T})(\bm{\tilde{S}}_{adj})^{l}||=||\bm{\tilde{U}}\text{diag}(\lambda_{1}^{l},...,\lambda_{N-1}^{l},0)\bm{\tilde{U}}^{T}||=\lambda_{N-1}^{l}$ (16)
where |||||$\cdot$||| | ⋅ | | denotes the spectral norm and the last equality follows from that the spectral norm of a diagonal matrix being equal to the largest absolute value of its diagonal entries. Thus, when $\lambda_{N-1}$ is close to 1, more layers are required to achieve the same level of convergence to the limiting state. Since $\lambda_{N-1}<1$, taking powers of $\lambda_{N-1}$ will eventually converge to zero, but the rate of this convergence slows dramatically as $\lambda_{N-1}$ approaches 1. Hence, because graphs with a large diameter and low maximum degree have a lower algebraic connectivity due to reduced inter-component connectivity, GNNs operating on such graphs will be less susceptible to over-smoothing when processing node features.
Nevertheless, our analysis demonstrates that large diameters and sparse connectivity generally mitigate over-smoothing. This insight motivates our proposal of benchmark datasets with significantly larger diameters than those currently used in the literature, refer back to Table 1. The underlying hypothesis is that reduced over-smoothing in high-diameter, sparse networks enables the possibility of GNNs learning representations that capture long-range dependencies when necessary.

### v2

@@ 3 Benchmarks: From Classical GNNs to Graph Transformers
In this section, we benchmark classical GNNs and GTs at different numbers of message-passing layers on City-Networks, and then contrast their behaviors with results on other commonly used graphs to examine the long-range dependencies across datasets.
Experimental setups. We consider transductive node classification with train/validation/test splits of $10\%/10\%/80\%$ on all City-Networks, in which we evaluate various classical GNNs and GTs. Specifically, we benchmark GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCNII (Chen et al., 2020) for GNNs; and GraphGPS (Rampášek et al., 2022), Exphormer (Shirzad et al., 2023), and SGFormer (Wu et al., 2023) for GTs. In addition, an MLP is also included to reflect the importance of graph topology (or lack thereof) in our task. Lastly, we fix the data split for benchmarking purposes and report the mean and standard deviation of the test accuracy over 5 runs with different random seeds.
Figure 3: Baseline results across datasets at different number of layers $L=[2,4,8,16]$. The results for GraphGPS are not shown on London as it is Out-of-Memory on our 4848GB GPU; the result for SGFormer on PascalVOC-SP is also not reported as it’s not originally designed for inductive setting.
Evaluation protocols. As a hypothesis, when a graph model is predicting on a certain node $v$, it will benefit from communicating with $v$’s neighbors from distant hops if the task is long-ranged. Following this rationale, we evaluate baseline models at different numbers of layers $L=[2,4,8,16]$, with the expectation that a larger $L$ would lead to better performance after sufficient training. By comparison, we also evaluate the same baselines on four representative datasets from the literature: Cora, ogbn-arxiv, Amazon-Ratings, and PascalVOC-SP from LRGB, which are homophilous, large-scale, heterophilous, and long-range dependent, respectively. We then compare the behaviors of baseline models across different datasets while increasing the number of layers $L$.
For fair comparison purposes, we follow the latest GNN tuning technique from Luo et al. (2024) that considers residual connection, normalization, dropout, etc. Readers are also referred to Appendix C.2 for full details on our training and experimental setups.
Results. From Figure 3, we can observe a clear and consistent improvement in performance for all baselines on our City-Networks as their message-passing depth $L$ increases from $2$ to $16$. It suggests that, even with sufficient training, a shallow graph model cannot outperform its deeper counterpart if it does not communicate with sufficiently distant hops. This result also means that the gains from incorporating long-range information as $L$ increases outweigh other side effects of depth. By contrast, the performance of baseline models gradually degrades as $L$ increases on Cora and ogbn-arxiv due to the locality of these tasks and depth-related issues. To elaborate on this matter, we further provide a theoretical justification for our dataset in Section 5, where we link the rate of over-smoothing (a prominent issue associated with depth) to the algebraic connectivity and network diameter of the underlying graph. On the other hand, the baseline performance remains relatively unchanged with only a slight increase on Amazon-Ratings and PascalVOC-SP, which suggests that these tasks are relatively short-ranged compared to City-Networks. To examine this hypothesis, we therefore introduce a quantitative measure of long-range dependency in the next section and compare it across datasets and baselines.
For completeness, we further conducted an ablation study in Appendix C.3 that (i) tests baselines under different hidden_size; and (ii) fixes the model size at $L=16$, refrains the model from seeing beyond hop-$H$ at each node, then tests baselines at different $H$. The results are all consistent with our above findings. In addition, we also summarize the best baseline performance on City-Networks in Appendix C.4, and discuss the potential limitation of our experiments in Appendix E.
4 A Long-range Measurement Based on Jacobians
@@
In this section, we propose to quantify the long-range dependency of a task given a trained model, by evaluating how the influence score $I(v,u)$ between node-pairs  (Xu et al., 2018; Gasteiger et al., 2022) varies with distance. The influence score measures the sensitivity of a GNN layer at node $v$ to the input feature of node $u$ using the Jacobian: $I(v,u)=\sum_{i}\sum_{j}\left|\,\partial\bm{H}^{(L)}_{vi}\big/\partial\bm{X}_{uj}\,\right|$, where $\bm{H}^{(L)}_{vi}$ is the $i^{th}$ entry of node $v$’s embedding at layer $L$, and $\bm{X}_{uj}$ is the $j^{th}$ entry of the input feature vector for node $u$. Unless otherwise specified, we assume $L$ refers to the last logit layer.
@@
where $T_{h}(v)=I_{\text{sum}}(v,h)$ is the total influence from the $h^{th}$-hop neighbors of node $v$, $\rho(v,u)$ is the length of the shortest path between $v$ and $u$ (note that this is equivalent to the $h$-hop shell later discussed in Section 6), and $N$ is the number of nodes in the network. Also, we would like to highlight that when $h=0$, $T_{0}(v)=I(v,v)$ becomes the influence of the feature at node $v$ on its output. The average total influence quantifies how much, on average, the features of nodes that are $h$ hops away affect the output at the focal node. In other words, by summing the Jacobian‐based influence scores from all $h$-hop neighbors and then averaging over all nodes, it provides an expected measure of the cumulative effect that distant nodes have on each node as the focal node.
@@
where $H$ is the maximum number of hops to be considered. Intuitively, one can understand $R$ as measuring how far away the average unit of influence is, where graphs with long-range dependencies are expected to have higher proportional influence between more distant nodes. Note that while the proposed $R$ bears some similarity to the recent work (Bamberger et al., 2025), they do not consider per-hop influence $T_{h}(v)$ as a diagnostic tool of the dependency decay. Finally, we refer readers to Appendix D for a discussion on the computational cost of this measurement.
@@
Table 2: Average size of the influence-weighted receptive field $R$ across different datasets and models.
@@
Model Paris Shanghai L.A. London Cora ogbn-arxiv Amazon PascalVOC
GCN 4.86 5.55 5.36 6.09 2.56 1.34 1.92 3.38
GraphSAGE 4.92 5.73 5.44 5.97 2.37 1.40 1.80 2.99
GCNII 3.62 3.64 3.68 3.66 2.76 3.5 1.79 2.60
GraphGPS 8.18 7.88 7.92 OOM 2.65 OOM 6.86 7.14
Exphormer 5.71 7.06 7.15 7.90 2.84 2.80 1.43 1.42
SGFormer 3.75 4.25 4.03 4.01 3.21 1.21 2.46 NA
Results. We validate $R$ and $\bar{T}_{h}$ on City-Networks using the baselines at $L=16$ layers from Section 3. Since PascalVOC-SP is under an inductive setting, we randomly sample 100 graphs from its testing set (more than $400k$ nodes in total) and report their average. The results in Table 2 show that $R$ is consistently higher on our City-Networks across all models compared to those on the other graph datasets, indicating a longer effective range of influence.
@@
Figure 4: Normalized average total influence $\bar{T}_{h}/\bar{T}_{0}$ averaged across nodes at different distances. Note that the influence calculation for GraphGPS is Out-of-Memory on London and ogbn-arxiv.
To better compare $\bar{T}_{h}$ across models and datasets, we normalize it by $\bar{T}_{0}$ and report $\bar{T}_{h}/\bar{T}_{0}$ in Figure 4. For all models, we generally observe a rapid influence drop at distant hops on Cora and ogbn-arxiv, whereas the decay is noticeably much slower on City-Networks. The patterns on Amazon-Ratings and PascalVOC-SP generally fall between the above two cases with influence concentrated on the first few hops, which further corroborate our findings in Section 3. While $R$ and $\bar{T}_{h}$ are both model-dependent, the above results yield a consistent ordering across datasets: models trained on City-Networks are required to leverage information from more distant hops to perform well, compared to those trained on existing datasets. For the rest of the paper, we will first provide a theoretical justification of the proposed dataset by linking over-smoothing to algebraic connectivity and network diameter in Section 5, and then present the intuition behind the proposed long-range measurement in Section 6.
5 Theoretical Justification of Topologies in City-Networks
@@
In this section, we provide a spectral analysis of over-smoothing in GNNs to justify our dataset design. As such, our main goal is not an in-depth theoretical analysis, since both over-smoothing and spectral analysis of GNNs have been carried out extensively in the literature (Rusch et al., 2023; Wu et al., 2023; Oono & Suzuki, 2020; Rong et al., 2019); instead, our aim is to build on top of these and derive simple results which provide a theoretical grounding of our dataset construction. The section is organized as follows. First, we review the concept of information loss in the limit of infinite GNN layers due to over-smoothing, which effectively vanishes the original node features and makes representations collapse to a value that is only dependent on the graph topology (Section 5.1 and Appendix F.1). Next, we relate the rate at which over-smoothing happens and the eigenvalues of the normalized adjacency operator in linearized GCNs (Section 5.2). Given that eigenvalues with larger absolute values slow down over-smoothing, we analyze how the magnitude of the eigenvalues relates to the graph topology of our datasets. This provides justification that sparse graphs with large diameters promote bigger positive eigenvalues for the normalized adjacency operator and hence are less vulnerable to over-smoothing (Theorem 5 and Appendix F.3). Finally, we argue that less over-smoothing makes it possible for GNNs to capture long-range dependencies in the case of our proposed dataset. All proofs are presented in Appendix F.2.
@@ 5.1 Preliminaries: A Spectral Perspective on Over-smoothing for Linear GCNs
We begin by analyzing over-smoothing through the lens of linearization, following the framework established in Wu et al. (2019) to build Simple Graph Convolutional Networks (SGCs). We select this model because it is a linearized and more interpretable version of the widely adopted GCN model, which has also been used for theoretical analysis in previous works such as (Giovanni et al., 2022).
@@
Let $\mathcal{G}=(V,E)$ be an undirected graph with $N=|V|$ nodes, adjacency matrix $\bm{A}\in\mathbb{R}^{N\times N}$ and degree matrix $\bm{D}\in\mathbb{R}^{N\times N}$ ($\bm{D}_{ii}=\sum_{j}\bm{A}_{ij}$). The normalized (augmented) adjacency operator $\bm{\tilde{S}}_{adj}\in\mathbb{R}^{N\times N}$ is defined as $\bm{\tilde{S}}_{adj}=(\bm{I}+\bm{D})^{-\frac{1}{2}}(\bm{I}+\bm{A})(\bm{I}+\bm{D})^{-\frac{1}{2}}$, where the original adjacency matrix and degree matrix have been augmented with self-loops. Wu et al. (2019) discuss the spectral properties of the normalized adjacency operator, which satisfies: $\lambda_{N}=1$, and $|\lambda_{i}|<1$ for all $i<N$. Therefore $\lim_{l\to\infty}(\bm{\tilde{S}}_{adj})^{l}=\bm{\tilde{u}}_{N}\bm{\tilde{u}}_{N}^{T}$, where $\bm{\tilde{u}}_{N}$ is the eigenvector corresponding to $\lambda_{N}$, whose entries at node $v$ are $\sqrt{1+\text{degree}(v)}$.
Via linearization (detailed in Appendix F.1), as the number of layers tends to infinity, the graph convolution operation collapses all node features to scalar multiples of $\bm{\tilde{u}}_{N}$, resulting in complete loss of the original feature information, that is, the learned representations suffer from over-smoothing.
5.2 Over-smoothing Rate on Algebraic Connectivity, Diameter, and Sparsity
@@
We now provide intuition about the types of graph topologies that would mitigate the over-smoothing problem. In particular, sparse graphs—those lacking the small-world effect commonly found in citation networks—and graphs with large diameters tend to experience over-smoothing at a slower rate. We argue that this slower rate of over-smoothing implies a higher likelihood that GNNs can learn long-range dependencies during optimization. We validate this intuition by bounding the eigenvalues of the normalized adjacency matrix using graph properties, then showing how these eigenvalues influence over-smoothing, and thus how controlling these graph properties can reduce the likelihood of over-smoothing. More concretely, we extend (Wu et al., 2019, Theorem 1), which only covers $\lambda_{1}$ and $\lambda_{N}$, to Proposition 4 which we prove in Appendix F.2, covering the normalized algebraic connectivity $\lambda_{N-1}$ and showing that $\lambda_{N-1}\left(\bm{S}_{adj}\right)<\lambda_{N-1}(\tilde{\bm{S}}_{adj})$.
@@
Assume a connected graph $\mathcal{G}$ with more than two nodes. For all $\gamma>0$,
@@
This allows us to relate the second largest positive eigenvalue of $\bm{\tilde{S}}_{adj}$ to the topology of graphs in the proposed dataset, using bounds similar to those in (Chung, 1997, Lemma 1.14).
@@
Dataset Paris Shanghai L.A. London Cora arxiv Ratings Pascal
Lower Bound 0.4741 0.6344 0.6095 0.5921 0.0324 -0.0639 0.1224 0.4857
From Table 3, we see that the lower bound in Equation (5) is generally higher for our datasets; this is because the bound is decreasing in $d_{max}$ and increasing in $diam(\mathcal{G})$, showing graphs with large diameter and low maximum degree will be more resilient to over-smoothing. We further provide the rationale behind this lower bound in Appendix F.3.